How far can a man drive out into the country at an average rate of 30 mph and return over the same road at an average of 50 mph, if he travels a total time of 8 hours?
Answer by checkley77(12844)

D=RT
D=30X DRIVE OUT
D=50(8-X) RETURN TRIP.
DISTANCES ARE EQUAL:
30X=50(8-X)
30X=400-50X
30X+50X=400
80X=400
X=400/80
X=5 HOURS FOR THE DRIVE OUT.
8-5=3 HOURS FOR THE RETURN TRIP.
PROOF:
30*5=50*3
150=150 MILES IS THE ONE WAY DISTANCE OR
2*150=300 MILES FOR THE ENTIRE RETURN TRIP.